# Making Every Head Count: Sparse Attention Without the Speed-Performance Trade-off

### v1

@@ Abstract
The design of Large Language Models (LLMs) has long been hampered by a fundamental conflict within their core attention mechanism: its remarkable expressivity is built upon a computational complexity of $O(H\cdot N^{2})$ that grows quadratically with the context size ($N$) and linearly with the number of heads ($H$). This standard implementation harbors significant computational redundancy, as all heads independently compute attention over the same sequence space. Existing sparse methods, meanwhile, often trade information integrity for computational efficiency. To resolve this efficiency-performance trade-off, we propose SPAttention, whose core contribution is the introduction of a new paradigm we term Principled Structural Sparsity. SPAttention does not merely drop connections but instead reorganizes the computational task by partitioning the total attention workload into balanced, non-overlapping distance bands, assigning each head a unique segment. This approach transforms the multi-head attention mechanism from $H$ independent $O(N^{2})$ computations into a single, collaborative $O(N^{2})$ computation, fundamentally reducing complexity by a factor of $H$. The structured inductive bias compels functional specialization among heads, enabling a more efficient allocation of computational resources from redundant modeling to distinct dependencies across the entire sequence span. Extensive empirical validation on the OLMoE-1B-7B and 0.25B-1.75B model series demonstrates that while delivering an approximately two-fold increase in training throughput, its performance is on par with standard dense attention, even surpassing it on select key metrics, while consistently outperforming representative sparse attention methods including Longformer, Reformer, and BigBird across all evaluation metrics. Our work demonstrates that thoughtfully designed structural sparsity can serve as an effective inductive bias that simultaneously improves both computational efficiency and model performance, opening a new avenue for the architectural design of next-generation, high-performance LLMs.
@@
Datasetshttps://huggingface.co/datasets/mlfoundations/dclm-baseline-1.0
@@ Introduction
$\begin{split}\text{Complexity}_{\text{SPAttention}}&=\sum_{h=0}^{H-1}O(N\cdot W_{h})=O\left(N\cdot\sum_{h=0}^{H-1}W_{h}\right)\\ &=O(N\cdot N)=O(N^{2})\end{split}$ (7)
@@
We directly compared SPAttention with standard dense attention across two model scales. As presented in Table 1, on the OLMoE 0.25B-1.75B scale, SPAttention outperforms the gold-standard dense attention on four out of six evaluation metrics, with significantly improved overall average performance. This advantage is further amplified at the larger OLMoE 1B-7B scale, where SPAttention achieves substantial performance gains on five out of six metrics, with only a minor disadvantage on Winogrande, while still maintaining a significant lead in overall average performance over standard attention.
This core observation strongly validates the central hypothesis of our methodology: the over-flexibility of standard MHA leads to computational redundancy and sub-optimal learning. By introducing a principled structured inductive bias, SPAttention compels functional specialization among heads, liberating finite computational resources from redundant local modeling and precisely redirecting them to modeling dependencies at distinct distance spans. This a priori design effectively prevents the model from settling into inefficient local optima, thereby fostering more efficient and generalizable representations.
Regarding the few metrics where standard attention holds a slight edge, we do not see this as an intrinsic flaw. Rather, it reflects different learning dynamics under a finite training budget. With its unconstrained connections, standard attention can ”brute-force” fit the data distribution of certain tasks with its full, redundant capacity. In contrast, SPAttention’s structured design encourages the model to learn more fundamental, general-purpose patterns. We posit that these minor performance discrepancies will narrow or even reverse with further training. Holistically, achieving comparable or better performance at roughly half the computational cost unequivocally demonstrates the value of SPAttention as a non-compromise solution.
@@
The comparison involved four methods: (1) a Standard Attention baseline; (2) our complete SPAttention method; and (3) three ablation variants: Sliding Window, ExclusiveBands (EBALL), and GappedBands (GBHALF). The attention patterns of our ablation variants are visualized in Figure 2.
@@
These results demonstrate that SPAttention’s superior performance stems not from a single trick, but from the synergistic outcome of its core principles: completeness, exclusivity, and balance. Any deviation from this holistic design incurs a significant performance penalty, establishing SPAttention as a robust and powerful general-purpose architecture.
@@ Conclusion
Our empirical validation demonstrates that SPAttention performs on par with or better than standard dense attention and consistently outperforms leading sparse methods. Our work’s crucial insight is that thoughtfully designed structural sparsity serves as a powerful inductive bias, simultaneously improving computational efficiency and model capabilities by eradicating redundancy at its source. This paves a promising path for designing next-generation models like LLaMA (Touvron et al. 2023) and Mamba (Gu and Dao 2023) while mitigating their expensive pre-training costs.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62572389, 72293581, 72274152, 62402376). We also acknowledge the support from Xi’an Jiaotong University.

### v2

@@ Abstract
The design of Large Language Models (LLMs) has long been hampered by a fundamental conflict within their core attention mechanism: its remarkable expressivity is built upon a computational complexity of O(H·N²) that grows quadratically with the context size (N) and linearly with the number of heads (H). This standard implementation harbors significant computational redundancy, as all heads independently compute attention over the same sequence space. Existing sparse methods, meanwhile, often trade information integrity for computational efficiency. To resolve this efficiency-performance trade-off, we propose SPAttention, whose core contribution is the introduction of a new paradigm we term Principled Structural Sparsity. SPAttention does not merely drop connections but instead reorganizes the computational task by partitioning the total attention workload into balanced, non-overlapping distance bands, assigning each head a unique segment. This approach transforms the multi-head attention mechanism from H independent O(N²) computations into a single, collaborative O(N²) computation, fundamentally reducing complexity by a factor of H. The structured inductive bias compels functional specialization among heads, enabling a more efficient allocation of computational resources from redundant modeling to distinct dependencies across the entire sequence span. Extensive empirical validation on the OLMoE-1B-7B and 0.25B-1.75B model series demonstrates that while delivering an approximately two-fold increase in training throughput, its performance is on par with standard dense attention, even surpassing it on select key metrics, while consistently outperforming representative sparse attention methods including Longformer, Reformer, and BigBird across all evaluation metrics. Our work demonstrates that thoughtfully designed structural sparsity can serve as an effective inductive bias that simultaneously improves both computational efficiency and model performance, opening a new avenue for the architectural design of next-generation, high-performance LLMs.
@@ Introduction
$\text{Complexity}_{\text{SPA}}$ $=\sum_{h=0}^{H-1}O(N\cdot W_{h})=O\left(N\cdot\sum_{h=0}^{H-1}W_{h}\right)$
$=O(N\cdot N)=O(N^{2})$ (7)
@@
We directly compared SPAttention with standard dense attention across two model scales. As presented in Table 1, on the OLMoE 0.25B-1.75B scale, SPAttention outperforms standard dense attention on four out of six evaluation metrics, with significantly improved overall average performance. This advantage is amplified at the larger OLMoE 1B-7B scale, where SPAttention achieves gains on five out of six metrics with only a minor disadvantage on Winogrande, maintaining a significant lead in overall average performance.
This strongly validates our hypothesis: the over-flexibility of standard MHA leads to redundancy. By introducing a principled structured inductive bias, SPAttention compels functional specialization, liberating resources from redundant local modeling to capture distinct dependencies. This design prevents the model from settling into inefficient local optima, fostering more generalizable representations.
Regarding the few metrics where standard attention holds a slight edge, we attribute this to its ability to “brute-force” fit specific distributions with unconstrained capacity, whereas SPAttention’s structured design encourages learning fundamental patterns. Holistically, achieving comparable or better performance at roughly half the computational cost unequivocally demonstrates the value of SPAttention as a non-compromise solution.
@@
The comparison involved four methods: (1) a Standard Attention baseline; (2) complete SPAttention method; and (3) three ablation variants: Sliding Window, ExclusiveBands (EBALL), and GappedBands (GBHALF). The attention patterns of our ablation variants are visualized in Figure 2.
@@
These results demonstrate that SPAttention’s superior performance stems not from a single trick, but from the synergistic outcome of its core principles: completeness, exclusivity, and balance. Any deviation from this holistic design incurs a significant performance penalty, establishing SPAttention as a robust general-purpose architecture.
@@ Conclusion
Our empirical validation demonstrates that SPAttention performs on par with or better than standard dense attention and consistently outperforms leading sparse methods. Our work’s crucial insight is that thoughtfully designed structural sparsity serves as a powerful inductive bias, simultaneously improving computational efficiency and model capabilities by eradicating redundancy at its source. This paves a promising path for pruning (Hu et al. 2025) and designing next-generation models like LLaMA (Touvron et al. 2023) and Mamba (Gu and Dao 2023) while mitigating their expensive pre-training costs.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62572389, 72293581, 72274152, 62402376).